# ul. Pasteura 5, 02-093 Warsaw, Poland

### v1

@@ 3 String Theory and the First Half of the Universe
As suggested towards the end of Section 2.1.1, fundamental strings with a time-varying tension typically arise when moduli fields acquire a non-trivial time dependence. In this section we give some concrete examples of alternative, early universe cosmologies arising in a String Theory setting, mostly inspired from Apers:2024dtn (see also Cicoli:2023opf for a review). The generic situation where exotic, stringy epochs are expected to appear is that of scalar fields (moduli) displaced from their minimum, which start rolling along (typically steep) potentials. In realistic scenarios, where other forms of energy density are present, this can lead to a cascade of events, resulting in the multi-phase cosmologies reviewed here, comprising kination, tracker and matter domination epochs. Let us also remark that these scenarios also assume inflation took place, although they are agnostic with respect to the actual mechanism. To be compatible with observational constraints, such non-standard epochs have to end before the beginning of BBN, and satisfy the constraints from $\Delta N_{\rm{eff}}$ Allahverdi:2020bys .
@@ 3.1 Compactifications at large volume
As remarked in Section 2.1.1, cosmic superstrings (in absence of warping) require a very large compactification volume in order to be phenomenologically viable. For this reason, we will consider explicit examples where moduli stabilization occurs at very large values of the volumes. The paradigmatic example is the Large Volume Scenario (LVS) Balasubramanian:2005zx ; Conlon:2005ki , arising from flux compactifications of Type IIB String Theory (See Conlon:2006gv for a review). It is one of the few known cases where all moduli can be stabilized, offering a rich and interesting phenomenology Conlon:2005jm ; Conlon:2006tq ; Conlon:2007gk ; Blumenhagen:2009gk ; Cicoli:2012sz ; Cicoli:2012aq ; Hebecker:2014gka . While such a specific choice might appear restrictive, the vast majority of what we are going to say does not depend on the fine details of the construction and equally applies to different vacua in the asymptotics of moduli space. All that is needed is a steep (close to exponential) potential in the canonically normalized volume that leads to a vacuum for exponentially large values of $\mathcal{V}\gtrsim 10^{10}$. The LVS is simply one of the best developed scenarios and allows one to make computations very concrete, for example by giving numerical values to $\mathcal{O}(1)$ coefficients. Another advantage is that most physical observables are (for the most part) determined by a single parameter, the volume $\mathcal{V}$. In Section 4, we will see how this leads to a unique prediction for the GW spectrum once the overall amplitude of the signal (equivalently, the string tension) are fixed, for example by experimental data. In particular, as we will discuss, natural volumes in LARGE volume scenarios give string tensions and reconnection probabilities in a range that could explain the PTA signal Ellis:2023tsl ; Avgoustidis:2025svu .
@@
$\Phi=\sqrt{\frac{2}{3}}M_{P}\log\mathcal{V},$ (22)
@@
$V_{LVS}(\Phi)=V_{0}\left[(1-\epsilon_{LVS}\Phi^{3/2})e^{-\lambda\frac{\Phi}{M_{P}}}+\delta e^{-\sqrt{6}\frac{\Phi}{M_{P}}}\right],$ (23)
where $\lambda=\sqrt{\frac{27}{2}}$ for LVS. It has a minimum for $\Phi\simeq 1/\epsilon_{LVS}^{2/3}$, which can be tuned in terms of microscopic parameters such as $g_{s}$, allowing the volume at the minimum to take exponentially large values. From a microscopic perspective, this results from the interplay between non-perturbative correction to the superpotential and the well known $\mathcal{N}=2$ BBHL correction Becker:2002nn ; Bonetti:2016dqh to the Kähler potential appearing at order $\mathcal{O}(\alpha^{\prime 3})$. Moroever, if one assumes the validity of the anti-D3 brane uplift, the coefficient $\delta$ can be fine-tuned to achieve a small, positive cosmological constant at the minimum.
As well as the LVS, one can consider more general mechanisms (potentials) relying on additional $\alpha^{\prime}$ and $g_{s}$ perturbative corrections to the large volume scalar potential (See Burgess:2020qsc ; Cicoli:2021rub for a systematic analysis). Examples include logarithmic field redefinitions involving the volume modulus Weissenbacher:2019mef ; Weissenbacher:2020cyf ; Klaewer:2020lfg or $\mathcal{N}=1$ mixed corrections to the Kähler potential at order $\mathcal{O}(g_{s}^{2}\alpha^{\prime 3})$ Antoniadis:1997eg ; Minasian:2015bxa ; Antoniadis:2018hqy ; Antoniadis:2019rkh ; Leontaris:2022rzj . In general, the tree-level Kähler potential is given by $K=-3\log\mathcal{V}$, where the volume $\mathcal{V}$ has a power-law dependence on the Kähler moduli $T_{i}$. It follows that any perturbative correction to the potential, once expressed terms of in canonically normalized fields, will consist of exponential terms. Likewise, logarithmic field redefinitions will manifest as power-law correction these exponential terms. For this reason, we can schematically parametrize the form of more general, large volume scalar potentials as
$V_{\rm{LV},1}(\Phi)=V_{0}\left[(1-\epsilon\Phi^{\kappa})e^{-\lambda_{1}\frac{\Phi}{M_{P}}}+\delta e^{-\lambda_{2}\frac{\Phi}{M_{P}}}\right],$ (24)
@@
$V_{\rm{LV},2}(\Phi)=V_{0}\left[Ae^{-\lambda_{1}\frac{\Phi}{M_{P}}}+Be^{-\lambda_{2}\frac{\Phi}{M_{P}}}+Ce^{-\lambda_{3}\frac{\Phi}{M_{P}}}\right]\,,$ (25)
which only features a minimum at large volume if there is a hierarchy between the coefficients. We discuss an example in Appendix A, inspired by a recent analysis of brane-antibrane inflation Cicoli:2024bwq . Here the constants $V_{0},\kappa,\lambda_{1},\lambda_{2},\lambda_{3}$, $\delta,A,B$ and $C$ depend on the case under consideration. From a qualitative level, the cosmological evolution outlined in the next section will not be affected significantly. However, it is clear that the different scenarios may give rise to quantitatively different predictions for the GW spectrum in Section 5.
@@ 3.1.1 Kination, trackers and all that
It is commonly assumed that, after inflation, the universe underwent a long epoch of radiation domination before finally reheating. However, this period is almost completely unconstrained by observations Allahverdi:2020bys , and alternatives are certainly on the table. In Conlon:2022pnx ; Apers:2024ffe , it was suggested that string theoretic considerations would indeed paint a very different picture, due to the evolution of the moduli fields. In the following, we will assume the presence of a single, cosmologically active modulus, the volume $\Phi$. This is justified by the existence of the LVS compactifications (and related scenarios) described in the previous section, where all other moduli can stabilized with parametrically higher masses.
@@
To reproduce the observed spectrum of primordial perturbations, cosmic inflation is expected to have taken place at very high energy scales, of the order $\Lambda_{\rm{inf}}\lesssim 10^{16}\,{\rm{GeV}}$ (corresponding to $H_{\rm{inf}}\lesssim 10^{13}\,{\rm{GeV}}$, with the upper limit set by Planck Planck:2018vyg ), although we are agnostic about its origin. [^19] In particular, we do not necessarily identify $\Phi$ with the inflaton. With the potential given by (23), if $\Phi$ found itself too close the minimum of the potential after inflation, this would imply a parametric suppression of the string scale (which controls all the scales of the theory) below that of inflation, raising the question of how one might have constructed a successful embedding of inflation into String Theory in the first place. It therefore makes sense to consider the case where, after inflation, the evolution begins with $\Phi\sim\mathcal{O}(1)M_{P}$, which subsequently starts rolling along the (approximately) exponential part of the potential, giving rise to an epoch of kinetic energy domination (kination). Since kinetic energy redshifts as $\rho_{K}\sim a(t)^{-6}$, faster than any other known fluid, this will be especially relevant right at the end of inflation, where other sources can be diluted and will not immediately dominate the energy budget of the universe. Explicit examples of inflation terminating with a kination phase for the volume modulus in string compactifications can be found in Conlon:2008cj ; Burgess:2022nbx , where the latter either acts acts as the inflaton or the waterfall field terminating inflation Linde:1993cn (see also Chen:2024roo ). Although the example we have presented here is for a specific (albeit well-motivated) construction, we expect similar behavior to be generic within string theoretic constructions, due to the (typically steep) exponential potentials that often appear for the moduli. For instance, scalar field mediated, stringy cosmologies that aim to match observational data on the known evolution of the universe often lead to a kination epoch when extrapolated back into the past Andriot:2024jsh ; Andriot:2024sif .
@@
Whenever a scalar field starts rolling along a steep potential, its potential energy will be quickly depleted to the advantage of the kinetic one, until the former becomes almost negligible. In absence of other sources, the only only obstacle to its motion will be from Hubble friction, in a regime known as kination. In this limit, the equations of motion for a generic scalar $\phi$ [^20] We will reserve the symbol $\Phi$ for the volume modulus, and $\phi$ for a generic scalar. reduce to
$\ddot{\phi}+3H\dot{\phi}=0\quad\quad\quad{\rm and}\quad\quad\quad 3H^{2}=\frac{1}{2}\dot{\phi}^{2},$ (26)
@@
$\phi=\phi_{0}+\sqrt{\frac{2}{3}}M_{P}\ln\left(\frac{t}{t_{0}}\right)\,,\quad\quad\quad a(t)=a_{0}\left(\frac{t}{t_{0}}\right)^{1/3}\,,\quad\quad\quad\quad H=\frac{1}{3t},$ (27)
@@
$\frac{M_{P}^{2}}{3t_{0}^{2}}\sim V\left(\phi_{0}\right)\sim\Lambda_{\mathrm{inf}}^{4},$ (28)
@@
$P_{\gamma}=(\gamma-1)\rho_{\gamma},\quad\quad\quad 0<\gamma<2,$ (29)
@@
$\Omega_{\gamma}\left(t_{0}\right)\equiv\frac{\rho_{\gamma}\left(t_{0}\right)}{3H_{\mathrm{inf}}^{2}M_{P}^{2}}=\frac{\rho_{\gamma}\left(t_{0}\right)}{\Lambda_{\mathrm{inf}}^{4}}=\varepsilon_{0}\ll 1,$ (30)
@@
$t_{\gamma}=t_{0}\,\varepsilon_{0}^{\frac{1}{\gamma-2}},$ (31)
defined as the time where the kinetic and fluid energy become equal. [^21] For a full derivation, including factors of $\mathcal{O}(1)$, see Apers:2024ffe . What happens afterwards? To answer this question, let us now specialize to the situation where $\phi$ can be identified with the volume modulus $\Phi$. Sufficiently far from the minimum, the potential (24) can be approximated as a simple exponential,
$V(\Phi)=V_{0}e^{-\frac{\lambda(\Phi-\Phi_{0})}{M_{P}}}.$ (32)
We also assume a fluid is also present, with equation of state as in (29). In this case, the equations of motion can be famously rephrased Wetterich:1987fm ; Copeland:1997et ; Ferreira:1997hj as the autonomous, dynamical system
[…]
@@
$x=\frac{\dot{\Phi}}{M_{P}}\frac{1}{\sqrt{6}H},\qquad y=\sqrt{\frac{V(\Phi)}{3}}\frac{1}{M_{P}H},$ (34)
and derivatives are with respect to $N=\ln a$. For suitable values of the parameters, the evolution of the system will necessarily converge to an attractor solution where all components redshift in the same way, and whose energy densities have fixed ratios. Such solutions are known as trackers, because both the kinetic and potential energy of the scalar field track or mimic the behaviour of the fluid component, and redshift as the latter would if it were dominating the energy density of the universe alone. In formulae, the scale factor and Hubble rate behave as
$a(t)=a(t_{t})\left(\frac{t}{t_{t}}\right)^{2/3\gamma}\,,\quad\quad\quad\quad H=\frac{2}{3\gamma}\frac{1}{t},$ (35)
@@
$\phi=\phi(t_{t})+\frac{2M_{P}}{\lambda}\ln\left(\frac{t}{t_{t}}\right),$ (36)
@@
$\Omega_{k}=x^{2}=\frac{3}{2}\frac{\gamma^{2}}{\lambda^{2}}\,,\quad\Omega_{p}=y^{2}=\frac{3(2-\gamma)\gamma}{2\lambda^{2}}\,,\quad\Omega_{\gamma}=1-x^{2}-y^{2}=1-\frac{3\gamma}{\lambda^{2}},$ (37)
where the subscripts $k$ and $p$ denote kinetic and potential energies respectively. In practice, realistic values of $\gamma$ are $\gamma=1$ and $\gamma=\frac{4}{3}$, corresponding to a matter or radiation tracker respectively. Possible sources in the early universe are a population of Primordial Black Holes (PBHs) for the former Conlon:2022pnx ; Apers:2024ffe (See for example Bhaumik:2022pil ; Bhaumik:2022zdd ) or decay products from the volume (such as the volume axion) in the latter case Conlon:2022pnx . The tracker solution will last until the volume reaches the minimum of the potential, and starts oscillating around it. From this moment onwards, the evolution will proceed with a phase of matter domination, followed by reheating. Notice that the tracker phase is actually crucial to ensure the field does not overcome the potential barrier to infinity Barreiro:1998aj ; Huey:2000jx ; Brustein:2004jp ; Barreiro:2005ua ; Conlon:2008cj ; Conlon:2022pnx , solving the so-called overshoot problem Brustein:1992nk .

### v2

@@ 3 String Theory and the First Half of the Universe
As suggested towards the end of Section 2.1.1, fundamental strings with a time-varying tension typically arise when moduli fields acquire a non-trivial time dependence. In this section we give some concrete examples of alternative, early universe cosmologies arising in a String Theory setting, mostly inspired from [105] (see also [37] for a review). In order to make our calculations fully explicit, we specialise to a particular corner of the string theory landscape, type IIB flux compactifications at large volume. The generic situation where exotic, stringy epochs are expected to appear is that of scalar fields (moduli) displaced from their minimum, which start rolling along (typically steep) potentials. In realistic scenarios, where other forms of energy density are present, this can lead to a cascade of events, resulting in the multi-phase cosmologies reviewed here, comprising kination, tracker and matter domination epochs. Let us also remark that these scenarios also assume inflation took place, although they are agnostic with respect to the actual mechanism. To be compatible with observational constraints, such non-standard epochs have to end before the beginning of BBN, and satisfy the constraints from $\Delta N_{\rm{eff}}$ [38].
@@ 3.1 Compactifications at large volume
As remarked in Section 2.1.1, cosmic superstrings (in absence of warping) require a very large compactification volume in order to be phenomenologically viable. For this reason, we will consider explicit examples where moduli stabilization occurs at very large values of the volumes. The paradigmatic example is the Large Volume Scenario (LVS) [9, 10], arising from flux compactifications of Type IIB String Theory (See [106] for a review). It is one of the few known cases where all moduli can be stabilized, offering a rich and interesting phenomenology [107, 108, 109, 110, 111, 112, 113]. While such a specific choice might appear restrictive, the vast majority of what we are going to say does not depend on the fine details of the construction and equally applies to different vacua in the asymptotics of moduli space. All that is needed is a steep (close to exponential) potential in the canonically normalized volume that leads to a vacuum for exponentially large values of $\mathcal{V}\gtrsim 10^{10}$. The LVS is simply one of the best developed scenarios and allows one to make computations very concrete, for example by giving numerical values to $\mathcal{O}(1)$ coefficients. Another advantage is that most physical observables are (for the most part) determined by a single parameter, the volume $\mathcal{V}$. In Section 4, we will see how this leads to a unique prediction for the GW spectrum once the overall amplitude of the signal (equivalently, the string tension) are fixed, for example by experimental data. In particular, as we will discuss, natural volumes in LARGE volume scenarios give string tensions and reconnection probabilities in a range that could explain the PTA signal [7, 19].
@@
$\Phi=\sqrt{\frac{2}{3}}M_{P}\log\mathcal{V},$ (3.1)
@@
$V_{LVS}(\Phi)=V_{0}\left[(1-\epsilon_{LVS}\Phi^{3/2})e^{-\lambda\frac{\Phi}{M_{P}}}+\delta e^{-\sqrt{6}\frac{\Phi}{M_{P}}}\right],$ (3.2)
where $\lambda=\sqrt{\frac{27}{2}}$ for LVS. It has a minimum for $\Phi\simeq 1/\epsilon_{LVS}^{2/3}$, which can be tuned in terms of microscopic parameters such as $g_{s}$, allowing the volume at the minimum to take exponentially large values. From a microscopic perspective, this results from the interplay between non-perturbative correction to the superpotential and the well known $\mathcal{N}=2$ BBHL correction [114, 115] to the Kähler potential appearing at order $\mathcal{O}(\alpha^{\prime 3})$. Moreover, if one assumes the validity of the anti-D3 brane uplift, the coefficient $\delta$ can be fine-tuned to achieve a small, positive cosmological constant at the minimum.
As well as the LVS, one can consider more general mechanisms (potentials) relying on additional $\alpha^{\prime}$ and $g_{s}$ perturbative corrections to the large volume scalar potential (See [116, 117] for a systematic analysis). Examples include logarithmic field redefinitions involving the volume modulus [118, 119, 120] or $\mathcal{N}=1$ mixed corrections to the Kähler potential at order $\mathcal{O}(g_{s}^{2}\alpha^{\prime 3})$ [121, 122, 123, 124, 125]. In general, the tree-level Kähler potential is given by $K=-3\log\mathcal{V}$, where the volume $\mathcal{V}$ has a power-law dependence on the Kähler moduli $T_{i}$. It follows that any perturbative correction to the potential, once expressed terms in terms of canonically normalized fields, will consist of exponential terms. Likewise, logarithmic field redefinitions will manifest as power-law correction these exponential terms. For this reason, we can schematically parametrize the form of more general, large volume scalar potentials as
$V_{\rm{LV},1}(\Phi)=V_{0}\left[(1-\epsilon\Phi^{\kappa})e^{-\lambda_{1}\frac{\Phi}{M_{P}}}+\delta e^{-\lambda_{2}\frac{\Phi}{M_{P}}}\right],$ (3.3)
@@
$V_{\rm{LV},2}(\Phi)=V_{0}\left[Ae^{-\lambda_{1}\frac{\Phi}{M_{P}}}+Be^{-\lambda_{2}\frac{\Phi}{M_{P}}}+e^{-\lambda_{3}\frac{\Phi}{M_{P}}}\right]\,$, (3.4)
which only features a minimum at large volume if there is a hierarchy between the coefficients. We discuss an example in Appendix A, inspired by a recent analysis of brane-antibrane inflation [39]. Here the constants $V_{0},\kappa,\lambda_{1},\lambda_{2},\lambda_{3}$, $\delta,A$ and $B$ depend on the case under consideration. [^19] A fully realistic model would arrange them so that a small and positive contribution to the observed dark energy is reproduced - a task beyond the scope of this article. From a qualitative level, the cosmological evolution outlined in the next section will not be affected significantly. However, it is clear that the different scenarios may give rise to quantitatively different predictions for the GW spectrum in Section 5.
@@ 3.1.1 Kination, trackers and all that
It is commonly assumed that, after inflation, the universe underwent a long epoch of radiation domination before finally reheating. However, this period is almost completely unconstrained by observations [38], and alternatives are certainly on the table. In [31, 34], it was suggested that string theoretic considerations would indeed paint a very different picture, due to the evolution of the moduli fields. In the following, we will assume the presence of a single, cosmologically active modulus, the volume $\Phi$. This is justified by the existence of the LVS compactifications (and related scenarios) described in the previous section, where all other moduli can stabilized with parametrically higher masses.
@@
To reproduce the observed spectrum of primordial perturbations, cosmic inflation is expected to have taken place at very high energy scales, of the order $\Lambda_{\rm{inf}}\lesssim 10^{16}\,{\rm{GeV}}$ (corresponding to $H_{\rm{inf}}\lesssim 10^{13}\,{\rm{GeV}}$, with the upper limit set by Planck [126]), although we are agnostic about its origin. [^20] In particular, we do not necessarily identify $\Phi$ with the inflaton. With the potential given by (3.2), if $\Phi$ found itself too close the minimum of the potential after inflation, this would imply a parametric suppression of the string scale (which controls all the scales of the theory) below that of inflation, raising the question of how one might have constructed a successful embedding of inflation into String Theory in the first place. It therefore makes sense to consider the case where, after inflation, the evolution begins with $\Phi\sim\mathcal{O}(1)M_{P}$, which subsequently starts rolling along the (approximately) exponential part of the potential, giving rise to an epoch of kinetic energy domination (kination). Since kinetic energy redshifts as $\rho_{K}\sim a(t)^{-6}$, faster than any other known fluid, this will be especially relevant right at the end of inflation, where other sources can be diluted and will not immediately dominate the energy budget of the universe. Explicit examples of inflation terminating with a kination phase for the volume modulus in string compactifications can be found in [127, 71], where the latter either acts as the inflaton or the waterfall field terminating inflation [128] (see also [129]). Although the example we have presented here is for a specific (albeit well-motivated) construction, we expect similar behavior to be generic within string theoretic constructions, due to the (typically steep) exponential potentials that often appear for the moduli. For instance, scalar field mediated, stringy cosmologies that aim to match observational data on the known evolution of the universe often lead to a kination epoch when extrapolated back into the past [130, 131].
@@
Whenever a scalar field starts rolling along a steep potential, its potential energy will be quickly depleted to the advantage of the kinetic one, until the former becomes almost negligible. In absence of other sources, the only obstacle to its motion will be from Hubble friction, in a regime known as kination. In this limit, the equations of motion for a generic scalar $\phi$, [^21] We will reserve the symbol $\Phi$ for the volume modulus, and $\phi$ for a generic scalar. reduce to
$\ddot{\phi}+3H\dot{\phi}=0\quad\quad\quad{\rm and}\quad\quad\quad 3H^{2}=\frac{1}{2}\dot{\phi}^{2},$ (3.5)
@@
$\phi=\phi_{0}+\sqrt{\frac{2}{3}}M_{P}\ln\left(\frac{t}{t_{0}}\right)\,,\quad\quad\quad a(t)=a_{0}\left(\frac{t}{t_{0}}\right)^{1/3}\,,\quad\quad\quad\quad H=\frac{1}{3t},$ (3.6)
@@
$\frac{M_{P}^{2}}{3t_{0}^{2}}\sim V\left(\phi_{0}\right)\sim\Lambda_{\mathrm{inf}}^{4},$ (3.7)
@@
$P_{\gamma}=(\gamma-1)\rho_{\gamma},\quad\quad\quad 0<\gamma<2,$ (3.8)
@@
$\Omega_{\gamma}\left(t_{0}\right)\equiv\frac{\rho_{\gamma}\left(t_{0}\right)}{3H_{\mathrm{inf}}^{2}M_{P}^{2}}=\frac{\rho_{\gamma}\left(t_{0}\right)}{\Lambda_{\mathrm{inf}}^{4}}=\varepsilon_{0}\ll 1,$ (3.9)
@@
$t_{\gamma}=t_{0}\,\varepsilon_{0}^{\frac{1}{\gamma-2}},$ (3.10)
defined as the time where the kinetic and fluid energy become equal. [^22] For a full derivation, including factors of $\mathcal{O}(1)$, see [34]. What happens afterwards? To answer this question, let us now specialize to the situation where $\phi$ can be identified with the volume modulus $\Phi$. Sufficiently far from the minimum, the potential (3.3) can be approximated as a simple exponential,
$V(\Phi)=V_{0}e^{-\frac{\lambda(\Phi-\Phi_{0})}{M_{P}}}.$ (3.11)
We also assume a fluid is also present, with equation of state as in (3.8). In this case, the equations of motion can be famously rephrased [132, 133, 134] as the autonomous, dynamical system
$\left\{\begin{aligned} x^{\prime}(N)&=-3x-\frac{V^{\prime}(\Phi)}{V(\Phi)}\sqrt{\frac{3}{2}}y^{2}+\frac{3}{2}x\left[2x^{2}+\gamma\left(1-x^{2}-y^{2}\right)\right]\\ y^{\prime}(N)&=\frac{V^{\prime}(\Phi)}{V(\Phi)}\sqrt{\frac{3}{2}}xy+\frac{3}{2}y\left[2x^{2}+\gamma\left(1-x^{2}-y^{2}\right)\right]\\ H^{\prime}(N)&=-\frac{3}{2}H\left(2x^{2}+\gamma\left(1-x^{2}-y^{2}\right)\right)\\ \Phi^{\prime}(N)&=\sqrt{6}x\end{aligned}\right$., (3.12)
@@
$x=\frac{\dot{\Phi}}{M_{P}}\frac{1}{\sqrt{6}H},\qquad y=\sqrt{\frac{V(\Phi)}{3}}\frac{1}{M_{P}H},$ (3.13)
and derivatives are with respect to $N=\ln a$. For suitable values of the parameters, the evolution of the system will necessarily converge to an attractor solution where all components redshift in the same way, and whose energy densities have fixed ratios. Such solutions are known as trackers, [^23] We follow e.g. [37] in calling this regime a tracker. This situation is also called a (tracking) scaling solution in the literature - a term we avoid in this context to avoid confusion with the scaling regime of the strings. because both the kinetic and potential energy of the scalar field track or mimic the behaviour of the fluid component, and redshift as the latter would if it were dominating the energy density of the universe alone. In formulae, the scale factor and Hubble rate behave as
$a(t)=a(t_{t})\left(\frac{t}{t_{t}}\right)^{2/3\gamma}\,,\quad\quad\quad\quad H=\frac{2}{3\gamma}\frac{1}{t},$ (3.14)
@@
$\phi=\phi(t_{t})+\frac{2M_{P}}{\lambda}\ln\left(\frac{t}{t_{t}}\right),$ (3.15)
@@
$\Omega_{k}=x^{2}=\frac{3}{2}\frac{\gamma^{2}}{\lambda^{2}}\,,\quad\Omega_{p}=y^{2}=\frac{3(2-\gamma)\gamma}{2\lambda^{2}}\,,\quad\Omega_{\gamma}=1-x^{2}-y^{2}=1-\frac{3\gamma}{\lambda^{2}},$ (3.16)
where the subscripts $k$ and $p$ denote kinetic and potential energies respectively. In practice, realistic values of $\gamma$ are $\gamma=1$ and $\gamma=\frac{4}{3}$, corresponding to a matter or radiation tracker respectively. Possible sources in the early universe are a population of Primordial Black Holes (PBHs) for the former [31, 34] (See for example [135, 136]) or decay products from the volume (such as the volume axion) in the latter case [31]. The tracker solution will last until the volume reaches the minimum of the potential, and starts oscillating around it. A point of concern, dubbed the overshoot problem [137], is that the field might not be trapped in the minimum and runaway to the boundary of field space. As discussed in [138, 139, 140, 141, 127, 31], the Hubble friction induced by the fluid driving the tracker period is sufficient to ensure that this is the case for the type of potentials at hand. The tracker phase is thus crucial to ensure the field does not overcome the potential barrier to infinity. From this moment onwards, as we now discuss, the evolution will proceed with a phase of matter domination, followed by reheating.